OpenStudy (anonymous):

integrate 188/(x^2)sqrt(16x^2-81)dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 188 }{ x^2\sqrt{16x^2-81} }dx\]

OpenStudy (anonymous):

okay so I got it to this point: \[47\int\limits_{}^{}\frac{ 1 }{ x^2\sqrt{x^2-(\frac{ 9 }{ 4 }})^2 }\]

OpenStudy (anonymous):

I'm confused about what to do next, please help :)

OpenStudy (anonymous):

whoops forgot the dx on the last integral